OpenStudy (anonymous):

Find the absolute maximum and minimum of f(x; y) = 2x^2- y^2 + 6y on the disk x^ 2 + y^ 2 <= 16.

OpenStudy (anonymous):

@wikiemol @RH

OpenStudy (anonymous):

@mathgeek!

OpenStudy (anonymous):

@iforgot

OpenStudy (anonymous):

@dmezzullo

OpenStudy (anonymous):

Hint: it's a continuous function on closed domain, so it obtains it's máximum on the boundaries

OpenStudy (anonymous):

so all i need to do is to find the max and min of f(x,Y)

OpenStudy (anonymous):

you know, that the max, min are obtained at x^ 2 + y^ 2 = 16. Notice the equality

OpenStudy (anonymous):

so you can express y as function of x, and viceversa

OpenStudy (anonymous):

for example: \(y=\pm\sqrt{16-x^2}\)

OpenStudy (anonymous):

plug this into the equation of f(x,y) and just find where the derivative \(f_x=0\) .

OpenStudy (anonymous):

so \(f(x,y) = f(x,y(x))=2x^2 -19+x^2 \pm6\sqrt{16-x^2}\)

OpenStudy (anonymous):

can i make a trig subsitution

OpenStudy (anonymous):

sry, where there is 19 should be 16

OpenStudy (anonymous):

ok thank

OpenStudy (anonymous):

can i let x= 4cos(u)

OpenStudy (anonymous):

yes, that's can be too

OpenStudy (anonymous):

\(32cos^2u-16sin^2u +24sinu\)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

i have \[12\cos ^{2}u +24\sin u-16\]

OpenStudy (anonymous):

\[12\sin ^{2}u-24\sin u-4\]

OpenStudy (anonymous):

\(2(16cos^2u)-16sin^2u+24sinu=32cos^2u-16sin^2u+24sinu\)

OpenStudy (anonymous):

which can be writen as: \(32-32sin^2u-16sin^2u+24sinu=-48sin^2u+24sinu+32\)

OpenStudy (anonymous):

i cant see what you did to get this line 32cos2u−16sin2u+24sinu

OpenStudy (anonymous):

just substitute x=4cosu, y=4sinu into the f(x; y) = 2x^2- y^2 + 6y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now find the derivative of this ,set it to 0 and find values of u that make that happen

OpenStudy (anonymous):

i got pi/2 and \[\sin ^{-1}(1/4)\]

OpenStudy (anonymous):

thats it